OpenStudy (anonymous):

find ∫_1^3〖4x^3 〗(x^4+6)^10 dx

OpenStudy (anonymous):

\[∫_1^3〖4x^3 〗(x^4+6)^{10}) dx\]

OpenStudy (anonymous):

ok, what methods do you know to solve integrals?

OpenStudy (lalaly):

\[\int\limits_{1}^{3}4x^3(x^4+6)^{10}dx\] let u=x^4 +6 du=4x^3dx so the integration becomes \[\large{\int{u^{10}du}}\]\[=\frac{u^{11}}{11}\] substituting the x back in u\[\int\limits{4x^3(x^4+6)^{10}dx} = \frac{(x^4+6)^{11}}{11}\]now just evaluate at the limits

OpenStudy (anonymous):

thanks

OpenStudy (lalaly):

=)